Three 45-Ω lightbulbs and three 75-Ω lightbulbs are connected in series. (a) what is the total resistance of the circuit? (b) what is their resistance if all six are wired in parallel? (c) Show the circuit illustration.


1
Expert's answer
2022-04-06T18:36:53-0400

"R_1=3\cdot(45+75)=360~\Omega,"

"R_2=\frac{45\cdot 75}{3\cdot(45+75)}=9.4~\Omega."